OpenStudy (anonymous):

7 cubedroot5- 4 sqrt 5

OpenStudy (solomonzelman):

\(\normalsize\color{red}{ 7\sqrt[3]{5} -4\sqrt[3]{5} }\) like this ?

OpenStudy (solomonzelman):

or \(\normalsize\color{red}{ 7\sqrt[3]{5} -4\sqrt{5} }\) ?

OpenStudy (anonymous):

3.025

OpenStudy (anonymous):

yes the second one!

OpenStudy (anonymous):

and these are binomial radical expression @SolomonZelman

OpenStudy (solomonzelman):

I would simplify the first one, but I can't do anything to the second one -:(

OpenStudy (anonymous):

nevermind it cannot be simplified

OpenStudy (anonymous):

i have a different one @SolomonZelman

OpenStudy (anonymous):

3x sqrtxy+4xsqrtxy

OpenStudy (solomonzelman):

Yes, sure

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x\sqrt{xy} -4x\sqrt{xy} }\) like this ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x\sqrt{xy} -4x\sqrt{xy}~=~(3x-4x)\sqrt{xy}~=~-x\sqrt{xy} }\)

OpenStudy (anonymous):

thanks so much!

OpenStudy (solomonzelman):

anytime

OpenStudy (anonymous):

may i ask you another please? @SolomonZelman